# On an entropy inequality for quadratic forms and applications

### v2

@@ Abstract
We introduce Frostman conditions for bivariate random variables and study discretized entropy sum–product phenomena in both independent and dependent regimes.
In the independent setting, we show that for any non-degenerate rational quadratic form $\phi(x,y)$ and $0<s<1$, there exists some positive constant $\epsilon=\epsilon(\phi,s)$ such that any independent marginally $(s,C;s,C)$-Frostman random variable $(X,Y)$ satisfies
$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon)$
for sufficiently large $n$. The proof introduces a novel multi‑step entropy framework, combining the submodularity formula, the discretized entropy Balog-Szemerédi-Gowers theorem, and state-of-the-art results on the Falconer distance problem, to reduce general forms to a diagonal core case.
In the dependent setting, we obtain analogous bounds for special polynomials. As an application, we derive that there exists $\epsilon=\epsilon(s)>0$ such that, for any $\delta$-separated set $A\subset[0,1]$ of cardinality $\delta^{-s}$ satisfying some non-concentration condition and any subset $G\subseteq A\times A$ with $|G|\geq\delta^{\epsilon}|A|^{2}$, the $\delta$-covering numbers satisfy
$\mathcal{E}_{\delta}(A+_{G}A)+\mathcal{E}_{\delta}(\alpha\cdot A^{\wedge 2}+_{G}\beta\cdot A^{\wedge 2})\gg_{\alpha,\beta,s}C^{-O_{s}(1)}\delta^{-\epsilon}|A|,$
whenever $\delta$ is small enough.
@@ 1 Introduction
More precisely, we are interested in bounding the following quantity $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ from below, where $\phi(x,y)=a_{1}x^{2}+a_{2}xy+a_{3}y^{2}$ is a non-degenerate rational quadratic form, i.e. $a_{2}^{2}-4a_{1}a_{3}\neq 0$, $X$ and $Y$ are independent $(s,C)$-Frostman random variables whose ranges are bounded, and $s\in(0,1)$.
Throughout this paper, the quadratic polynomial $\phi(x,y)$ is called special if there exist univariate real-analytic functions $h,a,b$ such that $\phi(x,y)=h\bigl{(}a(x)+b(y)\bigr{)}$. A direct computation shows that a special quadratic polynomial is one of the following two forms: $\phi(x,y)=\alpha x^{2}+\beta y^{2}$ and $\phi(x,y)=\gamma xy$, for $\alpha,\beta,\gamma\in\mathbb{R}^{*}$.
@@
These definitions are not equivalent. At a common exponent level $(s_{1},s_{2})$, we have the implications
@@
Our first theorem addresses the case that $X$ and $Y$ are independent real random variables, $(X,Y)$ forms marginally $(s_{1},C_{1};S_{2},C_{2})$-Frostman, and $\phi(x,y)$ is a non-degenerate rational quadratic form.
Theorem 1.2.
Let $\phi(x,y)$ be a non-degenerate rational quadratic form, $0<s<1$ and $C_{1},C_{2},c>0$. There exist a positive constant $\epsilon=\epsilon(\phi,s)$ and an integer $N=N(\phi,s,C_{1},C_{2},c)$ such that the following statement holds: Let $X$ and $Y$ be independent real random variables whose ranges are contained in [c,c][-c,c]. Suppose that $(X,Y)$ is marginally $(s_{1},C_{1};S_{2},C_{2})$-Frostman, with $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then
$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon)$ (2)
@@
To show the sharpness of this theorem, we summarize the main statements of several examples in Section 7:
@@
    Any $(s,C)$-Frostman random variable $Z$ has the property that $H_{n}(Z)\geq ns-\log C+O(1)$ (see Example 7.1 for an example matching this bound). The lower bound in (2) says that the discretized entropies of $X+Y$ and $\phi(X,Y)$ cannot be small simultaneously. We show in Examples 7.1 and 7.2 that at least one of the terms $H_{n}(X+Y)$ or $H_{n}(\phi(X,Y))$ can be as small as $(1+o(1))ns$.
@@
We pause here to add several further remarks. Our proof of Theorem 1.2 presents a connection between $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ and the entropy of the “distance” function, i.e. $H_{n}\big{(}(X_{1}-Y_{1})^{2}+(X_{2}-Y_{2})^{2}\big{)}$, where $(X_{1},Y_{1})$, $(X_{2},Y_{2})$ are i.i.d. copies of $(X,Y)$. When $a_{1}=a_{3}=0$ and $a_{2}=1$, i.e., $\phi(x,y)=xy$, Máthé and O’Regan [12] established a different type of connection by replacing the entropy of the “distance” function with that of the “slope” function. By combining this with a generalization of Marstrand’s theorem due to Orponen, Shmerkin, and Wang in [21], they were able to prove Theorem 1.2 for the full range of $s$ for this specific bilinear form function. We do not see how their method could be adapted to the form of $\phi$ with $a_{1},a_{3}\neq 0$. Moreover, in terms of applications, i.e. sum-product type problems over $\mathbb{R}$ or $\mathbb{F}_{p}$, the two functions $xy$ and $a_{1}x^{2}+a_{2}xy+a_{3}y^{2}$ appear to exhibit distinct behaviors. It would be interesting to explore more connections between $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ and the entropy of other functions.
It is worth noting that the proof of Theorem 1.2 can be extended to more general settings. In this version, we present two results for dependent random variables and special polynomials. Further results will be included in a forthcoming update.
Theorem 1.3.
Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$ and $C_{1},C_{2},c>0$. There exist a positive constant $\epsilon=\epsilon(s)$ and an integer $N=N(\alpha,\beta,s,C_{1},C_{2},c)$ such that the following statement holds: Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$. Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then
$\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}\geq n(s+\epsilon),$
and
$\max\left\{\frac{H_{n}(X+Y)+H_{n}(X-Y)}{2},\,H_{n}(XY)\right\}\geq n(s+\epsilon)$
for all $n>N$.
Both Theorems 1.2 and 1.3 are consequences of the following theorem, which only requires a jointly Frostman condition.
Theorem 1.4.
Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$, $C\geq 1$ and $c>0$. There exist a positive constant $\epsilon=\epsilon(s)$ and an integer $N=N(\alpha,\beta,s,C,c)$ such that the following statement holds: For any bivariate real random variable $(X,Y)$ that is jointly $(s_{1},s_{2},C)$-Frostman and has range contained in $c,c]^{2}$, where $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$, we have
for all $n>N$.
Example 7.6 will illustrate that the first term $\frac{H_{n}(X)+H_{n}(Y)}{2}$ in above formula cannot be removed, and cannot be replaced by $\frac{1}{2}H(X,Y)$ as in Theorem 1.3.
@@
Covering number: non-degenerate quadratic polynomials.
The next theorem presents an application of Theorem 1.2. Here, by $\underline{\dim}_{B}$ and $H^{s}$, we mean the lower box dimension and the $s$-Hausdorff measure, respectively.
Theorem 1.5.
Let $\phi(x,y)$ be a non-degenerate rational quadratic form. For $0<s<1$, there exists a positive constant $\epsilon=\epsilon(s{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},\phi})$ such that for every Borel set $A\subset{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}[0,1]}$ with $H^{s}(A)>0$, we have
$\max\left\{\underline{\dim}_{B}(A+A),\overline{\dim}_{B}(\phi(A,A))\right\}\geq s+\epsilon,$
and
$\max\left\{\overline{\dim}_{B}(A+A),\underline{\dim}_{B}(\phi(A,A))\right\}\geq s+\epsilon.$
For $A\subset\mathbb{R}$ and $\delta>0$, by $\mathcal{E}_{\delta}(A)$ we mean the $\delta$-covering number of $A$. The above theorem can be rewritten in the language of covering numbers.
Theorem 1.6.
Let $\phi(x,y)$ be a non-degenerate rational quadratic form. Given $0<s<1$ and $C\geq 1$. Then there exist positive constants $\epsilon=\epsilon(\phi,s)$ such that the following holds for all $\delta>0$. Let $A\subset[0,1]$ be any $\delta$-separated set of cardinality $\lceil\delta^{-s}\rceil$. Suppose that for all intervals $I$ of length $l$ at least $\delta$, we have the non-concentration estimate
$(A\cap I)\leq Cl^{s}(\#A).$
Then we have
$\mathcal{E}_{\delta}(A+A)+\mathcal{E}_{\delta}(\phi(A,A))\gg_{\phi,s}C^{-O_{\phi,s}(1)}\delta^{-\epsilon}(\#A).$
These two applications are of independent interest, particularly when $\phi(x,y)$ is a non-bilinear form, and might be considered as an application of the Falconer distance problem. If $\phi(x,y)=xy$, then stronger results have been obtained by using a variety of techniques, see [5, 11, 12, 17] and references therein for current progress. Given this line of research, it is reasonable to expect different proofs of Theorem 1.5 with other methods. In the prime field setting, such a result has been proved by using tools from incidence geometry, see [9, Theorem 1.5] and [15, Theorem 1.3], for example. When $\phi(x,y)=x^{2}+y^{2}$, after posting this paper to Arxiv, Tuomas Orponen pointed out that one can use his proof of [22, Corollary 1.3] and recent results in [1] or [23] to obtain a quantitative bound for this special case.
Fractional covering number: special quadratic polynomials.
@@
$A+_{G}A=\{x+y:\,(x,y)\in G\},\qquad\alpha\cdot A^{\wedge 2}+_{G}\beta\cdot A^{\wedge 2}=\{\alpha x^{2}+\beta y^{2}:\,(x,y)\in G\}.$
@@
Then, for all $G\subset A\times A$ with $|G|\geq\delta^{\epsilon}|A|^{2}$, then
@@
Sketch the main ideas in the proof of Theorem 1.2.
@@ 2.1 Notations
Let $(\Omega,\mathfrak{B},P)$ be a probability space. A random variable $X$ is a measurable function from $(\Omega,\mathfrak{B},P)$ to another measuable space $(\Omega^{\prime},\mathfrak{B}^{\prime})$, on which a pushforward probability measure $\mu$ is induced, i.e., $\mu(B^{\prime})=P(X\in B^{\prime})$ for any $B^{\prime}\in\mathfrak{B}^{\prime}$. We will say that $\mu$ is the associated distribution measure of $X$. We denote the range of $X$ by $range(X)$, which collects all the possible values that $X$ can take. In this paper, all the measurable spaces will be taken to be subspaces of ${\mathbb{R}}^{m}$, equipped with the Borel $\sigma$-algebra or its sub-$\sigma$-algebra, and the Euclidean norm $|\cdot|$. The dimension $m$ is always regarded as an absolute constant. The range of any random variable is assumed to be a bounded set. When two random variables $X$ and $Y$ have the same distribution, we write $X\equiv Y$.

### v3

# Frostman random variables, entropy inequalities, and applications

@@ Abstract
We introduce Frostman conditions for bivariate random variables and study discretized entropy sum–product phenomena in both independent and dependent regimes. We prove that for any non-degenerate rational quadratic form $\phi(x,y)$ and parameter $0<s<1$, there exists a positive constant $\epsilon=\epsilon(\phi,s)$ such that
$\max\{H_{n}(X+Y),H_{n}(\phi(X,Y))\}\geq n(s+\epsilon)$
for sufficiently large $n$, where the precise conditions on $(X,Y)$ depend on the Frostman level. The proof introduces a novel multi‑step entropy framework, combining the submodularity formula, the discretized entropy Balog-Szemerédi-Gowers theorem, and state-of-the-art results on the Falconer distance problem, to reduce general forms to a diagonal core case.
As an application, we derive a result on a discretized sum-product type problem. In particular, for a $\delta$-separated set $A\subset[0,1]$ of cardinality $\delta^{-s}$, satisfying some non-concentration conditions, and a dense subset $G\subset A\times A$, there exists $\epsilon=\epsilon(s,\phi)>0$ such that
$\mathcal{E}_{\delta}(A+_{G}A)+\mathcal{E}_{\delta}(\phi_{G}(A,A))\gg\delta^{-\epsilon}(\#A)$
for all $\delta$ small enough. Here by $\mathcal{E}_{\delta}(A)$ we mean the $\delta$-covering number of $A$, $\{A+_{G}A:=\{x+y\colon(x,y)\in G\}$, and $\phi_{G}(A):=\{\phi(x,y)\colon(x,y)\in G\}$.
@@ 1 Introduction
More precisely, we are interested in bounding the following quantity $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ from below, where $\phi(x,y)=a_{1}x^{2}+a_{2}xy+a_{3}y^{2}$ is a non-degenerate rational quadratic form, i.e. $a_{2}^{2}-4a_{1}a_{3}\neq 0$, and $(X,Y)$ is a bivariate Frostman random variable with bounded range.
@@
These definitions are not equivalent. At a common exponent level $(s_{1},s_{2})$, they form a hierarchy:
@@
Remarkably, different levels of this hierarchy yield entropy inequalities with different strengths and applications.
Our results demonstrate how the Frostman hierarchy affects entropy bounds. We call a quadratic form diagonal if $\phi(x,y)=\alpha x^{2}+\beta y^{2}$, and the main results read as follows.
Theorem 1.2 (Jointly Frostman, diagonal forms).
Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$, $C\geq 1$ and $c>0$. There exist a positive constant $\epsilon=\epsilon(s)$ and an integer $N=N(\alpha,\beta,s,C,c)$ such that the following statement holds. For any bivariate real random variable $(X,Y)$ that is jointly $(s_{1},s_{2},C)$-Frostman and has range contained in $c,c]^{2}$, where $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$, we have
@@
Theorem 1.3 (Conditionally Frostman, general forms).
Let $\phi(x,y)$ be a non-degenerate rational quadratic form, $0<s<1$ and $C_{1},C_{2},c>0$. There exist positive constants $\epsilon=\epsilon(s)$, $\epsilon^{\prime}=\epsilon^{\prime}(\phi,s)$ and an integer $N=N(\phi,s,C_{1},C_{2},c)$ such that the following statement holds. Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$. Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then, for all $n>N$,
$\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq\begin{cases}n(s+\epsilon),\quad&\text{if }\phi(x,y)\text{ is diagonal},\\ n(s+\epsilon^{\prime}),&\text{if }\phi(x,y)\text{ is non-diagonal}.\end{cases}$
@@
Theorem 1.4 (Marginally Frostman with independence).
Let $\phi(x,y)$ be a non-degenerate rational quadratic form, $0<s<1$ and $C_{1},C_{2},c>0$. There exist positive constants $\epsilon=\epsilon(s)$, $\epsilon^{\prime}=\epsilon^{\prime}(\phi,s)$ and an integer $N=N(\phi,s,C_{1},C_{2},c)$ such that the following statement holds. Let $X$ and $Y$ be independent real random variables whose ranges are contained in [c,c][-c,c]. Suppose that $(X,Y)$ is marginally $(s_{1},C_{1};S_{2},C_{2})$-Frostman, with $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then, for all $n>N$,
$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq\begin{cases}n(s+\epsilon),\quad&\text{if }\phi(x,y)\text{ is diagonal},\\ n(s+\epsilon^{\prime}),&\text{if }\phi(x,y)\text{ is non-diagonal}.\end{cases}$ (2)
To show the sharpness of these results, we summarize the main statements of several examples in Section 7:
@@
    Any $(s,C)$-Frostman random variable $Z$ has the property that $H_{n}(Z)\geq ns-\log C+O(1)$ (see Example 7.1 for an example matching this bound). We show in Examples 7.1 and 7.2 that there are i.i.d. random variables $X$ and $Y$ such that at least one of the terms $H_{n}(X+Y)$ or $H_{n}(\phi(X,Y))$ can be as small as $(1+o(1))ns$.
@@
  4. 4.
    In Theorems 1.2 and 1.3, the necessity of different terms in these maxima (for example, $\frac{H_{n}(X,Y)}{2}$ or $\frac{H_{n}(X)+H_{n}(Y)}{2}$) reflects subtle aspects of how dependency affects entropy, we provide an example showing it cannot be simplified in the Statement of Theorem 1.2 (Example 7.5).
@@
  7. 7.
    Theorem 1.2 can be extended to a more general form of
    $\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(f(X)+g(Y))\right\},$
    where $f$ and $g$ are polynomials with $\deg(f)=\deg(g)\geq 2$. Establishing such a result would require extending most of the arguments in this paper, and we plan to provide further updates in a subsequent version.
We pause here to add several further remarks. Our proof of Theorems 1.21.4 presents a connection between $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ and the entropy of the “distance” function, i.e. $H_{n}\big{(}(X_{1}-Y_{1})^{2}+(X_{2}-Y_{2})^{2}\big{)}$, where $(X_{1},Y_{1})$, $(X_{2},Y_{2})$ are i.i.d. copies of $(X,Y)$. When $a_{1}=a_{3}=0$ and $a_{2}=1$, i.e. $\phi(x,y)=xy$, Máthé and O’Regan [12] established a different type of connection by replacing the entropy of the “distance” function with that of the “slope” function. By combining this with a generalization of Marstrand’s theorem due to Orponen, Shmerkin, and Wang in [21], they were able to prove Theorem 1.4 for the full range of $s$ for this specific bilinear form function. We do not see how their method could be adapted to the form of $\phi$ with $a_{1},a_{3}\neq 0$. Moreover, in terms of applications, i.e. sum-product type problems over $\mathbb{R}$ or $\mathbb{F}_{p}$, the two functions $xy$ and $a_{1}x^{2}+a_{2}xy+a_{3}y^{2}$ appear to exhibit distinct behaviors. It would be interesting to explore more connections between $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ and the entropy of other functions.
@@
For $A\subset\mathbb{R}$ and $\delta>0$, by $\mathcal{E}_{\delta}(A)$ we mean the $\delta$-covering number of $A$.
@@
$A+_{G}A=\{x+y:\,(x,y)\in G\},\qquad\phi_{G}(A,A)=\{\phi(x,y):\,(x,y)\in G\},$
$\alpha\cdot A^{\wedge 2}+_{G}\beta\cdot A^{\wedge 2}=\{\alpha x^{2}+\beta y^{2}:\,(x,y)\in G\}.$
@@
Then, for all $G\subseteq A\times A$ with $G\geq\delta^{\epsilon}(\#A)^{2}$, we have
@@
Thanks to Theorem 1.3, extend this result to all non-degenerate rational quadratic forms is possible, but requiring minimum degree constraints.
Theorem 1.6.
Let $\phi(x,y)$ be a non-degenerate rational quadratic form, $0<s<1$ and $C\geq 1$. Then there exist positive constants $\epsilon=\epsilon(\phi,s)$ and $\delta_{0}=\delta_{0}(\alpha,\beta,s,C)$ such that the following holds for all $0<\delta\leq\delta_{0}$. Let $A\subset{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}[0,1]}$ be any $\delta$-separated set of cardinality $\lceil\delta^{-s}\rceil$. Suppose that for all intervals $I$ of length $l$ at least $\delta$, we have the non-concentration estimate
$(A\cap I)\leq Cl^{s}(\#A).$
Then, for any subset $G\subseteq A\times A$ such that
$\{y:\,(x,y)\in G\}\geq\delta^{\epsilon}(\#A)\quad\text{for any}\quad x\in A,$
$\{x:\,(x,y)\in G\}\geq\delta^{\epsilon}(\#A)\quad\text{for any}\quad y\in A,$
we have
$\mathcal{E}_{\delta}(A+_{G}A)+\mathcal{E}_{\delta}(\phi_{G}(A,A))\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A).$
In contrast to Theorem 1.6, the next result, an application of Theorem 1.4, holds for every scale $\delta>0$ and is stable under taking subsets: it applies to any $B\subset A$ with $B\geq\delta^{\varepsilon}\#A$.
Theorem 1.7.
Let $\phi(x,y)$ be a non-degenerate rational quadratic form, $0<s<1$ and $C\geq 1$. Then there exist positive constants $\epsilon=\epsilon(\phi,s)$ such that the following holds for all $\delta>0$. Let $A\subset{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}[0,1]}$ be any $\delta$-separated set of cardinality $\lceil\delta^{-s}\rceil$. Suppose that for all intervals $I$ of length $l$ at least $\delta$, we have the non-concentration estimate
$(A\cap I)\leq Cl^{s}(\#A).$
Then, for any subset $B\subseteq A$ satisfying $B\geq\delta^{\epsilon}(\#A)$, we have
$\mathcal{E}_{\delta}(B+B)+\mathcal{E}_{\delta}(\phi(B,B))\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A).$
These applications are of independent interest, particularly when $\phi(x,y)$ is a non-bilinear form, and might be viewed as consequences of the Falconer distance problem. If $\phi(x,y)=xy$, then stronger results have been obtained by using a variety of techniques, see [5, 11, 12, 17] and references therein for current progress. Given this line of research, it is reasonable to expect different proofs of Theorems 1.51.7 with other methods. In the prime field setting, such a result has been proved by using tools from incidence geometry, see [9, Theorem 1.5] and [15, Theorem 1.3], for example. When $\phi(x,y)=x^{2}+y^{2}$ ($\alpha=\beta=1$) and $G=A\times A$, after posting this paper to Arxiv, Tuomas Orponen pointed out that one can use his proof of [22, Corollary 1.3] and recent results in [1] or [23] to obtain a quantitative lower bound for the following estimate
$\mathcal{E}_{\delta}(A+A)+\mathcal{E}_{\delta}(A^{\wedge 2}+A^{\wedge 2})\gg_{\alpha,\beta,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A).$
We now sketch the main ideas of proofs in this paper.
@@ 2.1 Notations
Let $(\Omega,\mathfrak{B},P)$ be a probability space. A random variable $X$ is a measurable function from $(\Omega,\mathfrak{B},P)$ to another measuable space $(\Omega^{\prime},\mathfrak{B}^{\prime})$, on which a pushforward probability measure $\mu$ is induced, i.e., $\mu(B^{\prime})=P(X\in B^{\prime})$ for any $B^{\prime}\in\mathfrak{B}^{\prime}$. We will say that $\mu$ is the associated distribution measure of $X$. We denote the range of $X$ by $range(X)$, which collects all the possible values that $X$ can take. In this paper, all the measurable spaces will be taken to be subspaces of ${\mathbb{R}}^{m}$, equipped with the Borel $\sigma$-algebra or its sub-$\sigma$-algebra, the Euclidean norm $|\cdot|$, and the Lebesgue measure $\mathcal{L}^{m}(\cdot)$. The dimension $m$ is always regarded as an absolute constant. The range of any random variable is assumed to be a bounded set. When two random variables $X$ and $Y$ have the same distribution, we write $X\equiv Y$.